Metallic iron crystallizes in cubic lattice (pc, f
Answer #1

In general, Iron is a BBC cubit unit cell, and the calculus of it's atoms will prove it:

length unit cell = 287 pm * 1 cm / 1010 pm = 2.87x10-8 cm

Volume unit cell = (2.87x10-8)3 = 2.36x10-23 cm3

mass of Iron = 2.36x10-23 * 7.87 = 1.86x10-22 g

Atoms = 1.86x10-22 * 6.02x1023 atoms/mol = 112 g * atoms

If the molecular mass of IRon is 55.87 then:

number of atoms = 112 / 55.8 = 2.01 atoms

Hope this helps.